a) A plate with positive charge is 0.05 m higher than a plate with negative charge, and the electric field has an intensity of E=5x104N/C. Determine the work performed by the electric field E when a charge of +5x10-6C moves from the negative plate to the positive plate.

b) Find the electric potential at a point located at 0.05m of a charge of +8x10-6C. Find the potential energy of a charge of 3x10-9C that is located at that same point.



1
Expert's answer
2021-10-20T16:47:32-0400

Electric field


"E=\frac{V}{d}\\V=Ed=5\times10^4\times0.05=2500V"

Work

"W=qV"

Put value

"W=5\times10^{-6}\times2500=0.0125J"

Part(b)

"V=\frac{Kq}{r}\\V=\frac{9\times10^9\times8\times10^{-6}}{0.05}=144\times10^4V"

Potential energy

W=qV


"W=3\times10^{-9}\times144\times10^4=4.32\times10^{-3}J"